OpenStudy (anonymous):

What is the equation for the axis of symmetry of the quadratic function f(x) = 5x2 + 30x - 16?

OpenStudy (anonymous):

For quadratics, the axis symmetry equation is -b/2a. so in the above equation, a=5, b=30, c=-16 So plug and chug and you get an answer offfff? :)

OpenStudy (anonymous):

-30/2*5 = -30/10 = -3 so the axis of symmetry will be x = -3